OpenStudy (anonymous):

Equations with fractions; Show work! -7x + 3 = -15/2 Will Give Medals!!

OpenStudy (ivettef365):

first you can subtract 3 on both sides, do you have -7x +3 -3 = -15/2 -3

OpenStudy (anonymous):

X=-3/2 7x+3=15/2 -3 -3 7x=9/2

OpenStudy (ivettef365):

it should be -15/2

OpenStudy (anonymous):

How do you subtract three from a decimal!

OpenStudy (anonymous):

X=-3/2 7x+3=-15/2 ____-3___-3__ 7x=-10 1/2 7(-3/2)=10 1/2

OpenStudy (anonymous):

like this 3.5 -3.0 0.5

OpenStudy (anonymous):

You confused me! Sorry! I get the last comment but not the one before that, Why are you starting with x= -3/2

OpenStudy (anonymous):

because that is what x equals

OpenStudy (ivettef365):

I got positive 3/2

OpenStudy (anonymous):

o wow your right it is positve

OpenStudy (ivettef365):

that is all he needs, the answer is 3/2

OpenStudy (anonymous):

Thank you but can I see it worked out!

OpenStudy (ivettef365):

this is how you show your work: -7x + 3 = -15/2 -7x +3 -3 = -15/2 -3 15 3 -7x = - --- - ----- 2 1 -15-6 -7x = -------- 2 -7x = -21/2

OpenStudy (ivettef365):

now divide by -7 so you have 1 7 21 1 - ---- x - ----- x = - ---- x - ----- 7 1 2 7

OpenStudy (ivettef365):

so x = 21/14 simplify dividing by 7 and you get 3/2

OpenStudy (ivettef365):

that x after the first - 1/7 is to show you is multiplication, is not actually an x

OpenStudy (anonymous):

OK! Thank you